Use substitution to solve:
x=3y-2
2x+4y=16

this is what I got:
2(3y-2)+4y=16
6y+4y-4=16
10y-4=16
10y-4+4=16+4
10y/10 = 20/10
y=2

Answer by josmiceli(19441)

Looks right
x=3y-2
2x+4y=16
this is what I got:
2(3y-2)+4y=16
6y+4y-4=16
10y-4=16
10y-4+4=16+4
10y/10 = 20/10
y=2
then
x = 3*(2) - 2
x = 6 - 2 = 4
check in the other equation
2*(4) + 4*(2) = 16
8 + 8 = 16
checks